OpenStudy (anonymous):

for the function f(x) = x^2-3x-4, determine g(x) =f(2x)

jimthompson5910 (jim_thompson5910):

\[\Large f(x) = x^2-3x+4\] \[\Large f(2x) = (2x)^2-3(2x)+4\] Simplify from here

OpenStudy (anonymous):

oh okay thanks

jimthompson5910 (jim_thompson5910):

tell me what you get

OpenStudy (anonymous):

4x^2-6x+4

jimthompson5910 (jim_thompson5910):

you got it

OpenStudy (anonymous):

what if g(x) = 2f(2x)?

OpenStudy (anonymous):

would it be 8x^2-12x+8?

jimthompson5910 (jim_thompson5910):

Then you need to find \[\Large 2f(2x) = 2(4x^2-6x+4)\]

jimthompson5910 (jim_thompson5910):

you are correct

jimthompson5910 (jim_thompson5910):

\[\Large 2f(2x) = 8x^2-12x+8\]

OpenStudy (anonymous):

okay thanks

jimthompson5910 (jim_thompson5910):

yw